Find an equation of the line containing the points (3,5) and (9,9)

Answer by nyc_function(2741)

We first need to find the slope.

Let m = slope


m = (9 - 5)/(9 - 3)


m = 4/6


m = 2/3


The slope of the equation we are looking for is 2/3.


We now plug one of the points given (your choice which point to use) and the slope into the point-slope formula and solve for y.

y - y1 = m(x - x1)....This is the point-slope formula found in most math books.


I will use the point (3,5) but you can use the other point if you'd like.


y - 5 = 2/3(x - 3)


y - 5 = (2/3)x - 2


y = (2/3)x - 2 + 5


y = (2/3)x + 3


Understand?